OpenStudy (anonymous):

Express the solution using natural logarithm. 3e^(1+x)=2.

jhonyy9 (jhonyy9):

- so i think 3lne(1+x)=2lne 3(1+x)lne=2lne 3(1+x)=2 3+3x=2 3x=-1 x= -1/3

OpenStudy (anonymous):

i think jonny the process may be like ts.. ln{e^(1+x)}=ln(2/3) 1+x=ln(2/3)

OpenStudy (anonymous):

so x =ln(2/3)-1.. is it right jonny???

jhonyy9 (jhonyy9):

yes you think it right